If x and y are natural numbers, then how many solutions does the equation
x + y = 12 have?

Answer by Theo(13342)

it depends on whether you include 0 as part of the natural number set.

natural number can be either positive integers or non-negative integers.

your equation is x + y = 12.

assuming non-negative integers, then the possible solutions are:

0 + 12
1 + 11
2 + 10
3 + 9
4 + 8
5 + 7
6 + 6

assuming positive integers only, then the possible solutions are the same except that you would remove the 0 as one of the possible solutions.

positive integers DISABLED_event_only= 6 possible ways to add x + y = 12

non-negative integers DISABLED_event_only= 7 possible ways to add x + y = 12